TCS failure about 03:00 --> lasted about 45mins. Javier R. and Claudio A. managed to fix it.
Very productive night:
Several SNe fields observed. C3; z-filter (deep fields) X3; z-filter (deep fields) X3; gri-filters C1; grizY-filters C2; grizY-filters standard star fields observed at the beginning, middle and end of night. OBSTAC choose several SPT-E fields toward the end of the night. Post-night calibrations observed. DJJ
Very productive night:
Several SNe fields observed.
C3; z-filter (deep fields)
X3; z-filter (deep fields)
X3; gri-filters
C1; grizY-filters
C2; grizY-filters
standard star fields observed at the beginning, middle and end of night.
OBSTAC choose several SPT-E fields toward the end of the night.
Post-night calibrations observed.
DJJ